# DL-Chain: Scalable and Stable Blockchain Sharding with High Concurrency via Dual-Layer Consensus

### v1

@@ 5.2 Overhead Reduction
In this section, we focus on reducing overhead while ensuring system security. To prevent equivocation or transaction manipulation attacks, the strawman design in Section 4.2 necessitates each FC to maintain states of multiple PSs to verify raw transactions, resulting in considerable overhead. To address this issue, we leverage intra-shard consensus in each PS to ensure transaction validity, preventing manipulation attacks by ensuring $quorum\ size>\#\ of\ malicious\ nodes$. In this case, blocks containing invalid transactions cannot collect enough votes to pass consensus, as the honest nodes will not vote for them. However, a corrupted PS can still attack safety by equivocation (i.e., forking). We leverage FCs to resolve forking issues, detectable by examining the block header.
@@
Optimal Quorum Size. Combining the inequality that #$\#\ of\ honest\ nodes$ $\geq quorum\ size$ which prevents silent attack in Section 5.1 and the inequality that $quorum\ size>\#\ of\ malicious\ nodes$ which prevents manipulation attack above, we derive the optimal quorum size $m/2+1$ for a PS of size $m$. Consequently, we carefully adjust PS sizes to ensure the proportion of malicious nodes is less than $1/2$ with a high probability according to Section 6.1. This allows honest nodes in a PS to broadcast only the block header they voted for to the FC, who then checks signatures within the received block header and guarantees the finalized proposer block will not form forks.
@@
5.3 Cross-shard Transaction Handling
@@
We illustrate an example in Figure 3 that confirms the cross-shard transactions (transferred from accounts in $PS_{1}$ to accounts in $PS_{K+1}$) included in block $B$. Initially, $PS_{1}$ conducts intra-shard consensus on block $B$, including transactions whose payees are in $PS_{K+1}$, and waiting for the finalization of block $B$. Then, the honest voter of block $B$ must broadcast a receipt to convince $PS_{K+1}$ that the deduction operation is confirmed (finalized). The receipt includes a Merkle proof (refer to [15] for details) to ensure the existence of the batch of cross-shard transactions in block $B$, and the header of block $A$ to ensure block $B$ is finalized. Then, the destination $PS_{K+1}$ records the batched cross-shard transactions in block $C$ after verifying the receipt. Finally, the cross-shard transactions are confirmed when block $C$ is finalized by block $D$.
@@
Discussion. Our mechanism is inspired by existing relay-based cross-shard transaction processing schema (see Section 2.1), which has been proven for security and eventual atomicity. The difference lies in using secure FCs to safeguard the confirmation of cross-shard transactions.
@@
In our system, three points guarantee the security of cross-shard transactions. Firstly, the confirmation of the deduction operation is secure since we guarantee the security of FCs after rigorous theoretical derivations and proof, as shown in Section 6.1. Secondly, using the PS and FC identification for participants (derived from epoch randomness and nodes’ identities recorded in the beacon chain), nodes can verify whether blocks from other PSs and FCs have passed the consensus. Thirdly, malicious nodes can not manipulate the batched cross-shard transactions since they cannot forge a receipt’s Merkle proof.
@@
Besides, the following two points guarantee the eventual atomicity of cross-shard transactions. (I), at least one honest voter broadcasts each finalized proposer block to the destination PS since we have ensured that $quorum\ size>\#\ of\ malicious\ nodes$. (II), the cross-layer view change replaces malicious leaders, so a well-behaved leader will eventually pick transactions in receipts.
5.4 Pipelining Mechanism
@@
In DL-Chain, the efficiency of the Dual-Layer Consensus is crucial. A naive approach would be having PSs wait for their respective FC’s finalizer block before proposing new blocks, and vice versa, which leads to mutual blocking between layers. We employ a pipelined strategy to ensure efficient Dual-Layer Consensus operation. This strategy enables PSs to propose blocks optimistically without waiting for finalization, while FCs consecutively finalize multiple proposer blocks from each PS without delay. Firstly, we enable PSs to conduct block proposal optimistically. PS nodes can vote for multiple proposer blocks without waiting for finalization, temporarily maintaining states post-execution. Secondly, we make FCs to conduct block finalization without waiting. FCs can consecutively produce finalizer blocks, finalizing multiple proposer blocks of varying heights from each of their PSs. However, FCs should guarantee that all chained proposer blocks have passed intra-shard consensus. In this case, the accumulated proposer blocks in one PS can be finalized simultaneously.
Figure 4: Dual-Layer Consensus and block header design.
5.5 Summary of DL-Chain
@@
6 Security Analysis and Discussion
In this section, we begin by computing the failure probability of DL-Chain within each epoch. With negligible failure probability (i.e., ensuring the epoch security with extremely high probability), we then prove that the system is guaranteed to be secure as long as the proportion of malicious nodes in each PS is less than $1/2$ and that in each FC is less than $1/3$.
6.1 Epoch Security
To guarantee epoch security, we must ensure the system failure probability is below a certain threshold. This section aims to derive the failure probability of DL-Chain in each epoch so that we can adjust the system parameters to ensure a negligible system failure probability. We require that the fractions of malicious nodes are less than $1/3$ within each finalizer committee and less than $1/2$ within each proposer shard during each epoch. The overview of the calculation is as follows. Firstly, we calculate the failure probability of a finalizer committee. Specifically, it is divided into two cases. Case 1: It contains at least $1/3$ fraction of malicious nodes. Case 2: It is honest, but its proposer shards contain at least $1/2$ fraction of malicious nodes. Secondly, as in existing work, we calculate the union bound over all FCs to bound the failure probability of the whole system.
We calculate the failure probability for a finalizer committee now. The network size, the fraction of malicious in the network, and the size of finalizer committees are denoted as $N,f,n$, respectively. Let $X$ denote the random variable of the number of malicious nodes in a finalizer committee. We leverage the hypergeometric distribution, similar to existing sharding blockchains [16, 19, 36], to calculate the probability of a finalizer committee has $X=x$ malicious nodes:
$Pr[X=x]\\ =\frac{\binom{f\cdot N}{x}\binom{N-f\cdot N}{n-x}}{\binom{N}{n}}.$ (3)
@@ 6.1 Epoch Security
Like prior research [36, 19, 24], we also calculate the upper bound failure probability, assuming each shard’s failure probability is independent. We calculate the union bound over $K=n/m$ PSs, which results in the upper bound of the failure probability of an honest FC (according to case 2 that exists at least one PS containing at least $1/2$ malicious nodes).
@@
$Pr[System\ Failure]\\ \leq C\cdot Pr[FC\ Failure].$ (8)
Ensuring Negligible Failure Probability. DL-Chain, like most previous blockchain sharding works [21, 36, 16, 19], must ensure a negligible failure probability within each epoch to maintain security. Based on expressions 4 and 5, we need to adjust the FC size $n$ and the PS size $m$ to make sure there is a small $\varepsilon$ existing so that:
@@
If we achieve a negligible $\varepsilon$ such that the expression 9 holds, epoch security is ensured. The specific settings and the probabilities are shown in Section 7.3.
6.2 Security Analysis of Dual-Layer Consensus
In this section, we define and prove the security of our Dual-Layer Consensus, encompassing both safety and liveness aspects.
Since each node is assigned into one proposer shard and one finalizer committee simultaneously, let $\mathrm{C}_{t}^{P,\mathrm{n}}$ and $\mathrm{C}_{t}^{F,\mathrm{n}}$ denote the proposer chain and finalizer chain output by a full node $\mathrm{n}$ at time $t$. If the blocks, recorded from the genesis block onward, that constitute chain $\mathrm{C}_{t}^{\mathrm{n}}$ are a subset of the blocks that constitute chain $\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$, it is denoted as $\mathrm{C}_{t}^{\mathrm{n}}\preceq\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$. Additionally, if the blocks constituting chain $\mathrm{C}_{t}^{\mathrm{n}}$ are a proper subset of the blocks constituting chain $\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$, it is denoted as $\mathrm{C}_{t}^{\mathrm{n}}\prec\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$. The Dual-Layer Consensus is secure if the following two properties are satisfied.
@@ 6.2 Security Analysis of Dual-Layer Consensus
In Dual-Layer Consensus, considering a finalizer committee $FC_{i}$ ($1\leq i\leq C$) and one of its corresponding proposer shard $PS_{j}$ ($(i-1)K+1\leq j\leq iK$), which has $m_{j}$ nodes, and a fraction $f_{j}$ being malicious. assuming the network is a partial-synchronous network, the Dual-Layer Consensus protocol instantiated with finalizer committee $FC_{i}$ and proposer shard $PS_{j}$ satisfies
@@
We first prove the safety. Specifically, we first prove that finalized blocks are free from forks. Subsequently, we demonstrate the validity of transactions within the finalized blocks. Suppose the finalizer committee is safe. Then, without loss of generality, $\mathrm{C}_{t_{1}}^{F,\mathrm{n_{1}}}\preceq\mathrm{C}_{t_{2}}^{F,\mathrm{n_{2}}}$ for any two nodes $\mathrm{n_{1}}$ and $\mathrm{n_{2}}$ from this finalizer committee and times $t_{1}$ and $t_{2}$. Let $B_{t}^{P}$ represent the latest block of $PS_{j}$ finalized by $FC_{i}$ at time $t$. As detailed in Section 4.2, honest nodes within $FC_{i}$ play a crucial role in cross-layer verification. They ensure that, during the consensus process, the most recent finalized proposer block (e.g., $B_{t^{\prime}}^{P}$) remains free from conflicts with the blocks forming the chain that concludes with $B_{t}^{P}$ (i.e., $B_{t}^{P}\preceq B_{t^{\prime}}^{P}$ and $t<t^{\prime}$). And since blocks are linked by the collision-resistant hash function, the sequence of the finalized proposer block observed by $\mathrm{n_{1}}$ is a prefix of $\mathrm{n_{2}}$’s sequence. It implies that $\mathrm{C}_{t}^{P,\mathrm{n_{1}}}\preceq\mathrm{C}_{t^{\prime}}^{P,\mathrm{n_{2}}}$, concluding the safety proof. Besides, when the proposer shard is running FBFT [6] with $m_{j}$ distinct nodes and a quorum of $q_{j}=m_{j}/2+1$, each proposer block is verified by at least one honest node, thus even if the finalizer committee only verify the header of proposer block, the validity of the raw transaction is preserved.
@@
FBFT [6] has the same security guarantee with PBFT [12], which is stated as follows. PBFT satisfies safety and liveness for a group of $g$ nodes using $12g/3+$ 1 as a quorum when the fraction of byzantine nodes is less than $1/3$.
@@
The Dual-Layer Consensus protocol, instantiated with finalizer committee $FC_{i}$ running FBFT with $g_{i}$ nodes ($FC_{i}$ contains $f_{i}$ fraction of malicious nodes and uses $q_{i}=2g_{i}/3+1$ as a quorum), and proposer shard $PS_{j}$ running FBFT  with $m_{j}$ nodes ($PS_{j}$ contains $f_{j}$ fraction of malicious nodes and uses $q_{j}=m_{j}/2+1$ as a quorum), under partial-synchronous network, satisfies
@@ 6.3 Discussions
Temporary Stagnation. DL-Chain can prevent temporary stagnation, which arises from reshuffling and ledger migration of corrupted shards. The rationale behind this is that we guarantee the liveness of corrupted shards. Nevertheless, DL-Chain cannot prevent the delay of the rotation of a malicious leader. Fortunately, it only incurs a minor delay, as it does not involve cross-shard ledger migration. Besides, several studies have been conducted to mitigate the impact of malicious leaders on sharding, and these solutions can be applied to our system. For instance, RepChain [20] introduces a reputation mechanism to effectively reduce the possibility of a node with malicious behavior being elected as a leader.
@@
Additional Attacks. DL-Chain can also resist some other type of attacks. For instance, the eclipse attack can isolate the leader of PS and FC by controlling all their incoming and outgoing connections. Fortunately, our cross-layer view change mechanism can replace a PS leader, and the original FBFT comes with a view change mechanism to replace the FC leader. In addition, malicious nodes may collude to replace the honest leader. For two reasons, malicious nodes cannot replace a well-behaved, honest leader. Firstly, malicious nodes cannot obtain the honest leader’s private key to forge evidence of attacks. Additionally, they cannot gather quorum size of Complain messages because other honest nodes will not follow suit. The attackers can launch transaction censorship attacks to censor transactions [39] intentionally. Red Belly [14] proposes a leaderless BFT-typed consensus protocol. It prevents the transaction censorship attack conducted by a single leader via merging the micro-blocks proposed by multiple nodes into one complete block. Moreover, it can also prevent the eclipse attack on the specific leader. Most importantly, we can also use the above consensus protocol to resist these attacks since the intra-shard consensus protocol is an alternative component.
@@
Cross-shard Transaction. Besides supporting transfer transactions, DL-Chain can also exploit exiting methods [34, 19, 6] to handle complex smart contracts that span multiple proposer shards (which can also be in different FCs). The rationale is that DL-Chain provides safe finalization of proposer blocks to confirm each execution step, similar to the finalization of deduction operation for cross-shard transfer transactions (refer to Section 5.3). The optimization of complex smart contracts processing is orthogonal to our study and can be considered our future work.
@@
Although increasing the number of shards may result in more cross-shard communications, the enhanced concurrency achieved through smaller shards outweighs the potential performance impact. The main reason is that the size of a cross-shard transaction (typically receipts or metadata) is usually smaller than an intra-shard transaction [6, 34]. Moreover, DL-Chain only transmits headers for cross-layer verification.
7 Implementation and Evaluation
7.1 Implementation

### v2

@@ V-B Overhead Reduction
In this section, we focus on reducing overhead while ensuring system security. To prevent equivocation and transaction manipulation attacks, the strawman design in Section IV-B necessitates each FC to maintain states of multiple PSs to verify raw transactions, resulting in considerable overhead. To address the overhead issue, we first aim to leverage intra-shard consensus in each PS to ensure transaction validity, thus preventing manipulation attacks. To achieve this, we configure the size of each shard based on rigorous theoretical calculations (Section VI-A) to ensure that quorumsize>#ofmaliciousnodesquorum\ size>\#\ of\ malicious\ nodes with extremely high probability. In this case, blocks containing invalid transactions cannot collect enough votes to pass consensus, as the honest nodes will not vote for them. However, a corrupted PS can still attack safety by equivocation (i.e., forking). We then leverage FCs to resolve forking issues (detailed in Section IV-B), detectable by examining the block header.
@@
Optimal Quorum Size. Combining the inequality that #ofhonestnodes\#\ of\ honest\ nodes $\geq quorum\ size$ which prevents silent attack in Section V-A and the inequality that quorumsize>#ofmaliciousnodesquorum\ size>\#\ of\ malicious\ nodes which prevents manipulation attack above, we derive the optimal quorum size $m/2+1$ for a PS of size $m$. Consequently, we carefully adjust PS sizes to ensure the proportion of malicious nodes is less than $1/2$ with a high probability according to Section VI-A. This allows honest nodes in a PS to broadcast only the block header they voted for to the FC, who then checks signatures within the received block header and guarantees the finalized proposer block will not form forks.
@@
V-C Cross-shard Transaction Handling
@@
We illustrate an example in Figure 3 that confirms the cross-shard transactions (transferred from accounts in $PS_{1}$ to accounts in $PS_{K+1}$) included in block $B$. Initially, $PS_{1}$ conducts intra-shard consensus on block $B$, including transactions whose payees are in $PS_{K+1}$, and waiting for the finalization of block $B$. Then, the honest voter of block $B$ must broadcast a receipt to convince $PS_{K+1}$ that the deduction operation is confirmed (finalized). The receipt includes a Merkle proof (refer to [merkleTree] for details) to ensure the existence of the batch of cross-shard transactions in block $B$, and the header of block $A$ to ensure block $B$ is finalized. Then, the destination $PS_{K+1}$ records the batched cross-shard transactions in block $C$ after verifying the receipt. Finally, the cross-shard transactions are confirmed when block $C$ is finalized by block $D$.
@@
Discussion. Our mechanism is inspired by existing relay-based cross-shard transaction processing schema (see Section II-A), which has been proven for security and eventual atomicity. The difference lies in using secure FCs to safeguard the confirmation of cross-shard transactions.
@@
In our system, three points guarantee the security of cross-shard transactions. Firstly, the confirmation of the deduction operation is secure since we guarantee the security of FCs after rigorous theoretical derivations and proof, as shown in Section VI-A. Secondly, using the PS and FC identification for participants (derived from epoch randomness and nodes’ identities recorded in the beacon chain), nodes can verify whether blocks from other PSs and FCs have passed the consensus. Thirdly, malicious nodes can not manipulate the batched cross-shard transactions since they cannot forge a receipt’s Merkle proof.
@@
Besides, the following two points guarantee the eventual atomicity of cross-shard transactions. (I), at least one honest voter broadcasts each finalized proposer block to the destination PS since we have ensured that quorumsize>#ofmaliciousnodesquorum\ size>\#\ of\ malicious\ nodes. (II), the cross-layer view change replaces malicious leaders, so a well-behaved leader will eventually pick transactions in receipts.
V-D Pipelining Mechanism
@@
In StableShard, the efficiency of the DoLC protocol is crucial. A naive approach would be having PSs wait for their respective FC’s finalizer block before proposing new blocks, and vice versa, which leads to mutual blocking between layers. We employ a pipelined strategy to ensure efficient DoLC protocol operation. This strategy enables PSs to propose blocks optimistically without waiting for finalization, while FCs consecutively finalize multiple proposer blocks from each PS without delay. Firstly, we enable PSs to conduct block proposal optimistically. PS nodes can vote for multiple proposer blocks without waiting for finalization, temporarily maintaining states post-execution. Secondly, we make FCs to conduct block finalization without waiting. FCs can consecutively produce finalizer blocks, finalizing multiple proposer blocks of varying heights from each of their PSs. However, FCs should guarantee that all chained proposer blocks have passed intra-shard consensus. In this case, the accumulated proposer blocks in one PS can be finalized simultaneously.
Figure 4: DoLC protocol and block header design.
V-E Summary of StableShard
@@
VI Security Analysis and Discussion
In this section, we begin by computing the failure probability of StableShard within each epoch. With negligible failure probability (i.e., ensuring the epoch security with extremely high probability), we then prove that the system is guaranteed to be secure as long as the proportion of malicious nodes in each PS is less than $1/2$ and that in each FC is less than $1/3$.
VI-A Epoch Security
To guarantee epoch security, we must ensure the system failure probability is below a certain threshold. This section aims to derive the failure probability of StableShard in each epoch so that we can adjust the system parameters to ensure a negligible system failure probability. We require that the fractions of malicious nodes are less than $1/3$ within each finalizer committee and less than $1/2$ within each proposer shard during each epoch. The overview of the calculation is as follows. Firstly, we calculate the failure probability of a finalizer committee. Specifically, it is divided into two cases. Case 1: It contains at least $1/3$ fraction of malicious nodes. Case 2: It is honest, but its proposer shards contain at least $1/2$ fraction of malicious nodes. Secondly, as in existing work [rapidchain, cochain, li2022jenga, gearbox], we calculate the union bound over all FCs to bound the failure probability of the whole system.
We calculate the failure probability for a finalizer committee now. The network size, the fraction of malicious in the network, and the size of finalizer committees are denoted as $N,f,n$, respectively. Let $X$ denote the random variable of the number of malicious nodes in a finalizer committee. We leverage the hypergeometric distribution, similar to existing sharding blockchains [sgxSharding, pyramid, rapidchain], to calculate the probability of a finalizer committee has $X=x$ malicious nodes:
[…]
@@ VI-A Epoch Security
Like prior research [rapidchain, pyramid, cochain], we also calculate the upper bound failure probability, assuming each shard’s failure probability is independent. We calculate the union bound over $K=n/m$ PSs, which results in the upper bound of the failure probability of an honest FC (according to case 2 that exists at least one PS containing at least $1/2$ malicious nodes).
@@
[…]
Ensuring Negligible Failure Probability. StableShard, like most previous blockchain sharding works [omniledger, rapidchain, sgxSharding, pyramid], must ensure a negligible failure probability within each epoch to maintain security. Based on expressions 4 and 5, we need to adjust the FC size $n$ and the PS size $m$ to make sure there is a small $\varepsilon$ existing so that:
@@
If we achieve a negligible $\varepsilon$ such that the expression 9 holds, epoch security is ensured. The specific settings and the probabilities are shown in Section VII-C.
VI-B Security Analysis of DoLC Protocol
In this section, we define and prove the security of our DoLC protocol, encompassing both safety and liveness aspects.
Since each node is assigned into one proposer shard and one finalizer committee simultaneously, let $\mathrm{C}_{t}^{P,\mathrm{n}}$ and $\mathrm{C}_{t}^{F,\mathrm{n}}$ denote the proposer chain and finalizer chain output by a full node $\mathrm{n}$ at time $t$. If the blocks, recorded from the genesis block onward, that constitute chain $\mathrm{C}_{t}^{\mathrm{n}}$ are a subset of the blocks that constitute chain $\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$, it is denoted as $\mathrm{C}_{t}^{\mathrm{n}}\preceq\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$. Additionally, if the blocks constituting chain $\mathrm{C}_{t}^{\mathrm{n}}$ are a proper subset of the blocks constituting chain $\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$, it is denoted as $\mathrm{C}_{t}^{\mathrm{n}}\prec\mathrm{C}_{t^{\prime}}^{\mathrm{n^{\prime}}}$. The DoLC protocol is secure if the following two properties are satisfied.
@@ VI-B Security Analysis of DoLC Protocol
In DoLC protocol, considering a finalizer committee $FC_{i}$ ($1\leq i\leq C$) and one of its corresponding proposer shard $PS_{j}$ ($(i-1)K+1\leq j\leq iK$), which has $m_{j}$ nodes, and a fraction $f_{j}$ being malicious. assuming the network is a partial-synchronous network, the DoLC protocol instantiated with finalizer committee $FC_{i}$ and proposer shard $PS_{j}$ satisfies
@@
We first prove the safety. Specifically, we first prove that finalized blocks are free from forks. Subsequently, we demonstrate the validity of transactions within the finalized blocks. Suppose the finalizer committee is safe. Then, without loss of generality, $\mathrm{C}_{t_{1}}^{F,\mathrm{n_{1}}}\preceq\mathrm{C}_{t_{2}}^{F,\mathrm{n_{2}}}$ for any two nodes $\mathrm{n_{1}}$ and $\mathrm{n_{2}}$ from this finalizer committee and times $t_{1}$ and $t_{2}$. Let $B_{t}^{P}$ represent the latest block of $PS_{j}$ finalized by $FC_{i}$ at time $t$. As detailed in Section IV-B, honest nodes within $FC_{i}$ play a crucial role in cross-layer verification. They ensure that, during the consensus process, the most recent finalized proposer block (e.g., $B_{t^{\prime}}^{P}$) remains free from conflicts with the blocks forming the chain that concludes with $B_{t}^{P}$ (i.e., $B_{t}^{P}\preceq B_{t^{\prime}}^{P}$ and $t<t^{\prime}$). And since blocks are linked by the collision-resistant hash function, the sequence of the finalized proposer block observed by $\mathrm{n_{1}}$ is a prefix of $\mathrm{n_{2}}$’s sequence. It implies that $\mathrm{C}_{t}^{P,\mathrm{n_{1}}}\preceq\mathrm{C}_{t^{\prime}}^{P,\mathrm{n_{2}}}$, concluding the safety proof. Besides, when the proposer shard is running FBFT [Harmony] with $m_{j}$ distinct nodes and a quorum of $q_{j}=m_{j}/2+1$, each proposer block is verified by at least one honest node, thus even if the finalizer committee only verify the header of proposer block, the validity of the raw transaction is preserved.
@@
FBFT [Harmony] has the same security guarantee with PBFT [pbft], which is stated as follows. PBFT satisfies safety and liveness for a group of $g$ nodes using $2g/3+1$ as a quorum when the fraction of byzantine nodes is less than $1/3$.
@@
The DoLC protocol, instantiated with finalizer committee $FC_{i}$ running FBFT with $g_{i}$ nodes ($FC_{i}$ contains $f_{i}$ fraction of malicious nodes and uses $q_{i}=2g_{i}/3+1$ as a quorum), and proposer shard $PS_{j}$ running FBFT  with $m_{j}$ nodes ($PS_{j}$ contains $f_{j}$ fraction of malicious nodes and uses $q_{j}=m_{j}/2+1$ as a quorum), under partial-synchronous network, satisfies
@@ VI-C Discussions
Temporary Stagnation. StableShard can prevent temporary stagnation, which arises from reshuffling and ledger migration of corrupted shards. The rationale behind this is that we guarantee the liveness of corrupted shards. Nevertheless, StableShard cannot prevent the delay of the rotation of a malicious leader. Fortunately, it only incurs a minor delay, as it does not involve cross-shard ledger migration. Besides, several studies have been conducted to mitigate the impact of malicious leaders on sharding, and these solutions can be applied to our system. For instance, RepChain [repchain] introduces a reputation mechanism to effectively reduce the possibility of a node with malicious behavior being elected as a leader.
@@
Additional Attacks. StableShard can also resist some other type of attacks. For instance, the eclipse attack can isolate the leader of PS and FC by controlling all their incoming and outgoing connections. Fortunately, our cross-layer view change mechanism can replace a PS leader, and the original FBFT comes with a view change mechanism to replace the FC leader. In addition, malicious nodes may collude to replace the honest leader. For two reasons, malicious nodes cannot replace a well-behaved, honest leader. Firstly, malicious nodes cannot obtain the honest leader’s private key to forge evidence of attacks. Additionally, they cannot gather quorum size of Complain messages because other honest nodes will not follow suit. The attackers can launch transaction censorship attacks to censor transactions [sandwich] intentionally. Red Belly [redBelly] proposes a leaderless BFT-typed consensus protocol. It prevents the transaction censorship attack conducted by a single leader via merging the micro-blocks proposed by multiple nodes into one complete block. Moreover, it can also prevent the eclipse attack on the specific leader. Most importantly, we can also use the above consensus protocol to resist these attacks since the intra-shard consensus protocol is an alternative component.
@@
Cross-shard Transaction. Besides supporting transfer transactions, StableShard can also exploit exiting methods [monoxide, pyramid, Harmony] to handle complex smart contracts that span multiple proposer shards (which can also be in different FCs). The rationale is that StableShard provides safe finalization of proposer blocks to confirm each execution step, similar to the finalization of deduction operation for cross-shard transfer transactions (refer to Section V-C). The optimization of complex smart contracts processing is orthogonal to our study and can be considered our future work.
@@
Although increasing the number of shards may result in more cross-shard communications, the enhanced concurrency achieved through smaller shards outweighs the potential performance impact. The main reason is that the size of a cross-shard transaction (typically receipts or metadata) is usually smaller than an intra-shard transaction [Harmony, monoxide]. Moreover, StableShard only transmits headers for cross-layer verification.
VII Implementation and Evaluation
VII-A Implementation